OpenStudy (anonymous):

he value of y varies directly with x, and y = 1 when x = 3. Find y when x = 18.

OpenStudy (anonymous):

The*

OpenStudy (anonymous):

3*6=18

OpenStudy (anonymous):

1*6=6....y=6

OpenStudy (anonymous):

y varies directly with x then y=kx where k is the constant of proportionality y = 1 when x = 3 1=k*3 so k=1/3 so y= x/3 so when x=18 y=18/3 =6